OpenStudy (anonymous):

Solve the equation 1296 x = 6

OpenStudy (vishweshshrimali5):

I can help you

OpenStudy (vishweshshrimali5):

See

OpenStudy (anonymous):

x =0,00462963

OpenStudy (vishweshshrimali5):

1296x = 6 => x = 6/1296 = 1/216

OpenStudy (vishweshshrimali5):

1/216 = 0.00462963 Satisfied ?

OpenStudy (vishweshshrimali5):

Are you satisfied